OpenStudy (anonymous):

Help? Choose the ordered pair (x,y) that is a solution of the equation. -15x – 2y = 23

OpenStudy (anonymous):

There are infinitely many solutions. (-1,-4) works

OpenStudy (anonymous):

-15x - 2y = 23 -2y = 15x + 23 y = -(15/2)x - 23/2 plug in any x and get a y. For example, lets plug in x = 0: y = -(15/2)(0) - 23/2 y = -23/2 So (0, 23/2) is a point that works.

OpenStudy (anonymous):

I mean (0, -23/2)

OpenStudy (anonymous):

A. (-3,11) B. (2,-9) C. (3,-11) D. (-2,9)

OpenStudy (anonymous):

lets plug in -3: y = -(15/2)(-3) - 23/2 y = 45/2 - 23/2 y = 22/2 = 11

OpenStudy (anonymous):

option A

OpenStudy (anonymous):

or just plug the ordered pairs into your original equation and see which one works out.